OpenStudy (macaylam):

A quarterback throws a football toward a receiver from a height of 6 ft. The initial vertical velocity of the ball is 14 ft/s. At the same time that the ball is thrown, the receiver raises his hands to a height of 8 ft and jumps up with an initial vertical velocity of 10 ft/s. h=-16t^2+14+6 h=-16t^2+10+8 After 0.5 seconds, the receiver's hands and the ball will be at the same height. The receiver's hands and the ball are at the same height at ___ ft above the ground.

OpenStudy (retireed):

Those equation don't make sense to me. The 14 and the 10 should have a t variable associated with them, right?

OpenStudy (macaylam):

h=-162^2+14+6 h=-16t^2+10+8 these?

OpenStudy (retireed):

yeah

OpenStudy (macaylam):

can you help me figure it out?

OpenStudy (retireed):

h=-16t^2+14t+6 h=-16t^2+10t+8 I'm not sure where the -16t^t comes from, but I guess it is gravity. Plug in for t =0.5 into either or both equations and see what you get.

OpenStudy (macaylam):

I plugged in 0.5 and got h=24 and h=14...

OpenStudy (retireed):

Those are incorrect or at least not what I got using two calculators and two graphing functions. Write down your steps one by one so I can find your mistake. I will try to figure out those values on my end.

OpenStudy (macaylam):

I put it into wolfram alpha

OpenStudy (retireed):

Take a screen shot and attach it.

OpenStudy (retireed):

You are still using the wrong equations. These are the correct equations. NOTE the t behind the 14 and the t behind the 10. h=-16t^2+14t+6 h=-16t^2+10t+8 The speed of the ball 14 ft/sec is dependent on time t The receiver jumping up at 10ft/sec is dependent on time t

OpenStudy (macaylam):

so I plug 0.5 into those equations?

OpenStudy (retireed):

Do you think that would be a good way to proceed?